Early Payment discount decisions:

Prairie Manufacturing has four possible​ suppliers, all of which offer different credit terms. Except for the differences in credit​ terms, their products and services are virtually identical. The credit terms offered by these suppliers are shown in the following​ table:

SUPPLIER        CREDIT TERM

J                      3/5        NET 30 EOM

K                      4/30      NET 100 EOM

L                      2/15      NET 60 EOM

M                     2/10      NET 120 EOM

.

​ (Assume a​ 365-day year.)

A. Calculate the approximate cost of giving up the cash discount from each supplier.

B. If the firm needs​ short-term funds, which are currently available from its commercial bank at

99​%, and if each of the suppliers is viewed separately​, which, if​ any, of the​ suppliers' cash discounts should the firm give​ up?  

C. Now assume that the firm could stretch by 30 days its accounts payable​ (net period​ only) from supplier M. What​ impact, if​ any, would that have on your answer in part b relative to this​ supplier?
Answer #1

Part (a)

Please see the table below. The last column contains the answer. Please be guided by the second row that explains how each column has been calculated.

Supplier Credit Terms Cash discount Credit days Discount Period cost of giving up the cash discount
D CD DP D / (1 - D) x 365 / (CD - DP)
J 3/5 net 30 EOM 3% 30 5 45.15%
K 4/30 net 100 EOM 4% 100 30 21.73%
L 2/15 net 60 EOM 2% 60 15 16.55%
M 2/10 net 120 EOM 2% 120 10 6.77%

Part (b)

Decision criterion: If cost of giving up the cash discount < bank rate of 9.0%, then the discount should be given up.

This criterion is met in case of supplier M where cost of giving up the cash discount = 6.77% < 9%

Hence, the firm should give up the supplier M’s cash discount.

Part (c)

Supplier M's terms now become 2/10 net (120 + 30) EOM i.e. 2/10 net 150 EOM

Hence, cost of giving up the cash discount = D / (1 - D) x 365 / (CD - DP) = 2% / (1 - 2%) x 365 / (150 - 10) = 5.32%, which is still < bank rate = 9.0%

Hence, our answer in part (b) with respect to supplier M remains unchanged.